40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 96 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  12134   Wed May 25 11:51:40 2016 SteveUpdatesafetySURF 2016 safety
Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

Varun has received 40m specific basic safety training today.

  12133   Wed May 25 08:32:55 2016 SteveUpdateSUSlocal EQ 3.5m

Local EQ 3.5 mag  at 2:28 UTC May 24, 2016 Rancho Cucamonga, Ca.....no damage

 

Attachment 1: 3.5Cucam.png
3.5Cucam.png
Attachment 2: local3.5cucam.png
local3.5cucam.png
  12132   Wed May 25 02:54:09 2016 ericqUpdateGeneralOdds and ends

WFS locking point seemed degraded; I hand aligned and reset the WFS offsets as usual.

ITMX oplev recentered. While doing so, I noticed an ETMX excursion rear its head for the first time in a long while :crying

There was no active length control on ETMX, only OSEM damping + oplevs. Afterwards, its still moving around with only local damping on. I'm leaving the oplevs off for now.

  12131   Tue May 24 23:17:37 2016 ericqUpdateCOCFinesse modelling - mode overlap scans

I think you should use the current actual PRC & SRC cavity lengths as measured, as it would be simplest to simply replace the folding mirror optics without changing the macroscopic lengths / optic positions. (EDIT: Gautam rightly points out that we have to move things around regardless, since our current lengths include propagation through the folding mirror subtrates)

Moreover, the recycling cavity lengths you posted are not the right "ideal" lengths to use, as they do not account for the complex reflectivities of the sidebands off of the arm cavities (I have made this mistake myself). See this 40m wiki page for details.

In short, given our current modulation frequency, the ideal lengths to use would be:

  • Ideal arm length of 37.795 m
  • Ideal PRC length of 6.753 m
  • Ideal SRC length of 5.399 m

These are the lengths that the recycling cavity optics were positioned for (though we did not achieve them perfectly). If you do a finer PRC/SRC length scan around the DRFPMI resonance of your model, you would presumably see some undesired sideband splitting. 

  12130   Tue May 24 22:49:02 2016 gautamUpdateCOCFinesse modelling - mode overlap scans

Summary:

Having played around with a toy finesse model, I went about setting up a model in which the RC folding mirrors are not flipped. I then repeated the low-level tests detailed in the earlier elog, after which I ran a few spatial mode overlap analyses, the results of which are presented here. It remains to do a stability analysis.

Overview of model parameters (more details to follow):

  • PRC length = 6.7727m (chosen using l_{PRC} = (N+\frac{1}{2})\frac{c}{2f_1}, N=0 - I adjusted the position of the PRM to realize this length in the model, while leaving all the other vertex optics in the same positions as in elog 9590
  • SRC length = 5.4182 (chosen using l_{SRC} = M\frac{c}{2f_2} but not l_{SRC} = N\frac{c}{2f_1}, M and N being integers, for M=2 - as above, I adjusted the position of the SRM to realize this in the model, while leaving all other vertex optics in the same positions as in elog 9590. It remains to be verified if it is physically possible to realize these dimensions in vacuum without any beam clipping etc but I think it should be possible seeing as the PRM and SRM had to be moved by less than 2cm from their current positions..
  • For the losses, I used the most recent numbers we have where applicable, and put in generic 25ppm loss for all the folding mirrors/BS/AR surfaces of arm cavity mirrors/PRM/SRM. Arm round trip loss was equally distributed between ITMs and ETMs
  • Arm lengths used: L_X = 37.79m, L_Y = 37.81m
  • To set the "tunings" of the various mirrors, I played around with a few configurations to see where the various fields resonated - it turns out that for PRM, ITMX, ITMY, ETMX and ETMY, the "phase" in the .kat file can be set as 0. while that for the SRM can be set as 90. In the full L1/H1 interferometer .kat files, these are tuned even further to the (tenth?!) decimal place, but I think these values suffice for out purposes.

Results (general note: positive RoC in these plots mean a concave surface as seen by the beam):

  • Attachments #1, #2 and #3 reproduce the low-level tests performed earlier for this updated model - i.e. I look at the arm transmission with no PRM/SRM, circulating PRC power with no ETMs, and circulating SRC power with no ETMs. Everything looks consistent here... In Attachment #2, there is no legend, but the (almost overlapping) red and green lines are meant to denote the +f1 and +f2 sidebands.
  • Attachments #4 and #5 are a summary of the mode-overlap scans for the PRC and SRC. What I did was to vary the radius of curvature of the RC mirrors (finesse only allows you to vary Rcx and Rcy, so I varied both simultaneously) and calculate the mode overlap between the appropriate pairs of cavities (e.g. PRX and XARM) in the tangential and saggital planes. The take-away here is that there is ~5% mode-mismatch going from an RoC of 1000m to 300m. I've also indicated the sag corresponding to a given RoC - these are pretty tiny, I wonder if it is possible to realize a sag of 1um? I suppose it is given that I've regularly seen specs of surface roughness of lambda/10?
  • Attachment #6 shows the PRC gain (calculated as T_PRC * (transmitted arm power with PRM / transmitted arm power without PRM) as a function of the RoC of PR2 and PR3. As a sanity check, I repeated this calculation with lossless HR surfaces (but with nominal 25ppm losses for AR surfaces of ITMs, and BS etc), shown in Attachment #7. I think these make sense too...
  • Attachment #8 - in order to investigate possible mode mismatch between the arm modes due to different radii of curvature of the ETMs, I kept the ETMY RoC fixed at 57.6m and varied the ETMY RoC between 50m and 70m (here, I've plotted the mode matching efficiency as a function of the RoC of the ETM in the X and Y directions separately - the mode overlap is computed as \frac{1}{\sqrt{2}}(x^2 + y^2) where x and y denote the overlap in the tangential and saggital planes respectively. It would seem that we only lose at most a couple of percent even if the RoCs are mismatched by up to 10m...
  • Attachment #9 - .kat file and the various pykat scripts used to generate these plots...

Next step is to carry out a stability analysis...

Attachment 1: armTransmission.pdf
armTransmission.pdf
Attachment 2: prcFSR.pdf
prcFSR.pdf
Attachment 3: srcTransmission.pdf
srcTransmission.pdf
Attachment 4: modeMatchPRX.pdf
modeMatchPRX.pdf
Attachment 5: modeMatchSRX.pdf
modeMatchSRX.pdf
Attachment 6: PRCgainScan.pdf
PRCgainScan.pdf
Attachment 7: PRCgainLossless.pdf
PRCgainLossless.pdf
Attachment 8: armModeMatchScan.pdf
armModeMatchScan.pdf
Attachment 9: Finesse_files.zip
  12129   Tue May 24 17:55:17 2016 VarunUpdateElectronicsUsing Altium

Contacted Charles regarding use of Altium. Got to know that Altium is installed on cit40m iMac in Win7 on VirtualBox. Had to update Virtualbox to get it working. Altium now works for sometime, but then fails, saying that it is unlicensed.

  12128   Tue May 24 10:21:36 2016 ericqSummarySUSITMX Oplev loops

I did a quick measurement of the ITMX oplev loops, both pitch and yaw have about the same upper UGF as previous measurements with the previous laser; about 4 Hz. 

  12127   Mon May 23 17:47:51 2016 VarunUpdateGeneralSURF 2016

Tested the AGC today with LSC cavity transmission signal and error signal. Not in real time still.

Key to attachments:

cav_tr-eps-converted-to.pdf: LSC cavity transmission signal input

cav_tr_out-eps-converted-to.pdf: LSC cavity transmission signal, output of the AGC.

Attachment 1: cav_tr-eps-converted-to.pdf
cav_tr-eps-converted-to.pdf
Attachment 2: cav_tr_out-eps-converted-to.pdf
cav_tr_out-eps-converted-to.pdf
Attachment 3: err-eps-converted-to.pdf
err-eps-converted-to.pdf
Attachment 4: err_out-eps-converted-to.pdf
err_out-eps-converted-to.pdf
  12126   Mon May 23 15:51:32 2016 steveSummarySUSoplev laser summary updated

 

Quote:

 

Quote:

 

Quote:

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

      May  13, 2014             ETMX,  .............laser in place 90 d

      May  22, 2012             ETMY, 

     Oct.  7,  2013             ETMY,  LT  503 d  or  1.4 y............bad beam quality ?

     Aug. 8,  2014              ETMY,  .............laser in place   425 days  or  1.2 y

 

      Sept. 5, 2014              new 1103P, sn P893516  installed at SP table for aLIGO oplev use qualification

     

           May 23, 2016             ITMX dead laser sn P845648 replaced after 1062 days [2.9 yrs] by 1103P, sn P859884, with output output  2.6 mW, nicely round beam quality at 15 meters.

Attachment 1: oplSum.png
oplSum.png
  12125   Mon May 23 10:55:49 2016 steveSummarySUSITMX oplev laser replaced

 

      May 23, 2016             ITMX dead He/Ne laser sn P845648 replaced after 1062 days [2.9 yrs] by 1103P, sn P859884, with output  2.6 mW, nicely round beam quality at 15 meters.

                                                                                                                                                    Power just before viewport 1 mW,  returning light on qpd 154 microW =  7,500 counts

 

Attachment 1: ITMXoplev.png
ITMXoplev.png
  12124   Fri May 20 17:36:06 2016 gautamUpdateLSCNew stands for TransMon/Oplev QPDs

As we realized during the EX table switch, the transmitted beam height from the arm is not exactly 4" relative to the endtable, it is more like 4.75" at the X-end (yet to be investigated at the Y-end). As a result, the present configuration involves the steering optics immediately before the Oplev and TransMon QPDs sending the beam downwards at about 5 degrees. Although this isn't an extremely large angle, we would like to have things more level. For this purpose, Steve has ordered some Aluminium I-beams (1/2 " thick) which we can cut to size as we require. The idea is to have the QPD enclosures mounted on these beams and then clamped to the table. One concern was electrical isolation - but Steve thinks Delrin washers between the QPD enclosure and the mount will suffice. We will move ahead with getting these machined once I investigate the situation at the Y end as well.. The I beams should be here sometime next week...

  12123   Fri May 20 00:06:19 2016 VarunUpdateGeneralSURF 2016

I have written a basic version of AGC, and have done some tests with a data file. will do tests on whitening and agc today. Also, today I have to go to the SSN office. Hence will be late.

 

-Varun

Quote:

Finished writing the code on whitening. I have to still test it. uploaded on github noise cancellation repo. @eric could you give me some data of noise power spectral density for testing the code?

-Varun

Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

 

 

  12122   Thu May 19 16:29:20 2016 SteveUpdateendtable upgradeOptical layout almost complete

 

 

Attachment 1: ETMX_4x3_closed.jpg
ETMX_4x3_closed.jpg
Attachment 2: sealedETMXenclosure.jpg
sealedETMXenclosure.jpg
  12121   Wed May 18 17:42:52 2016 VarunUpdateGeneralSURF 2016

Finished writing the code on whitening. I have to still test it. uploaded on github noise cancellation repo. @eric could you give me some data of noise power spectral density for testing the code?

-Varun

Quote:

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

 

  12120   Wed May 18 01:10:22 2016 gautamUpdateCOCFinesse modelling

I've been working on putting together a Finesse model for the current 40m configuration. The idea was to see if I could reproduce a model that is in agreement with what we have been seeing during the recent DRFPMI locks. With Antonio and EricQs help, I've been making slow progress in my forays into Finesse and pyKat. Here is a summary of what I have so far.

  • Arm lengths were taken from some recent measurements done by yutaro and me 
  • Recycling cavity lengths were taken from Gabriele's elog 9590 - it is likely that the lengths I used have errors ~1cm - more on this later. Furthermore, I've tried to incorporate the flipped RC folding mirrors - the point being to see if I can recover, for example, a power recycling gain of ~7 which is what was observed for the recent DRFPMI locks.
  • I used Yutaro's most recent arm loss numbers, and distributed it equally between ITM and ETM for modeling purposes. 
  • For all other optics, I assumed a generic loss number of 25ppm for each surface

Having put together the .kat file (code attached, but this is probably useless, the new model with RC folding mirrors the right way will be what is relevant), I was able to recover a power recycling gain of ~7.5. The arm transmission at full lock also matches the expected value (125*80uW ~ 10mW) based on a recent measurement I did while putting the X endtable together. I also tuned the arm losses to see (qualitatively) that the power recycling gain tracked this curve by Yutaro. EricQ suggested I do a few more checks:

  1. Set PRM reflectivity to 0, scan ETMs and look at the transmission - attachment #1 suggests the linewidth is as we expect 
  2. Set ETM reflectivity to 0, scan PRM - attachment #2 suggests a Finesse of ~60  for the PRC which sounds about right
  3. Set ETM reflectivity to 0, scan SRM and verify that only the 55 MHz sidebands resonate - Attachment #3

Conclusion: It doesn't look like I've done anything crazy. So unless anyone thinks there are any further checks I should do on this "toy" model, I will start putting together the "correct" model - using RC folding mirrors that are oriented the right way, and using the "ideal" RC cavity lengths as detailed on this wiki page. The plan of action then is

  • Evaluating the mode-matching integrals between the PRC and the arm cavities as a function of the radius of curvature of PR2 and PR3
  • Same as above for the SRC
  • PRC gain as a function of RoC of folding mirrors
  • Mode overlap between the modes from the two arm cavities as a function of the RoC of the two ETMs (actually I guess we can fix RoC of ETMy and just vary RoC of ETMx).

Sidenote to self: It would be nice to consolidate the most recent cavity length measurements in one place sometime...

Attachment 1: arms.pdf
arms.pdf
Attachment 2: PRC.pdf
PRC.pdf
Attachment 3: SRC.pdf
SRC.pdf
Attachment 4: Finesse_model.zip
  12119   Tue May 17 14:46:51 2016 SteveUpdateVACRGA scan at day 595

We have good RGA scan now. There was no scan for 3 months.

Attachment 1: RGAscan595d.png
RGAscan595d.png
Attachment 2: pd78-560Hz-d600.png
pd78-560Hz-d600.png
  12118   Tue May 17 05:50:43 2016 Varun KelkarUpdateGeneralSURF 2016

Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.

-Varun

  12117   Sun May 15 19:48:08 2016 SteveUpdateVACrun out of N2

3-4 hrs ago we run out of nitrogen. We are back to Vacuum Normal

 

 

Attachment 1: noN2.png
noN2.png
  12116   Thu May 12 14:29:58 2016 gautamUpdateVACRGA back up and running

It looks like the hardware reset did the trick. Previously, I had just tried ssh-ing into c0rga and rebooting it. At the time, however, Steve and I noticed that the various LEDs on the RGA unit weren't on, as they are supposed to be in the nominal operating state. Today, Steve reported that all LEDs except the RS232 one were on today, so I just tried following the steps in this elog again, looks like things are back up and running. I'm attaching a plot of the scan generated using plotrgascan MATLAB script, it looks comparable to the plot in elog 11697, which if I remember right, was acceptable.

Unless there is some reason we want to keep this c0rga machine, I will recommission one of the spare Raspberry Pis lying around to interface with the RGA scanner when I get the time...

Quote:
Quote:

Our last RGA scan is from February 14, 2016  We had a power outage on the 15th

Gautom has not succeded  reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?

The c0rga computer was off, I turned it on via front panel button. After running RGAset.py, RGAlogger.py seems to run. However, there are error messages in the output of the plotrgascan MATLAB script; evidiently there are some negative/bogus values in the output. 

I'll look into it more tomorrow.

This is a cold scan.

Attachment 1: RGAscan_12May2016.png
RGAscan_12May2016.png
  12115   Wed May 11 16:39:01 2016 ericqUpdateVACc0rga alive, output wonky
Quote:

Our last RGA scan is from February 14, 2016  We had a power outage on the 15th

Gautom has not succeded  reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?

The c0rga computer was off, I turned it on via front panel button. After running RGAset.py, RGAlogger.py seems to run. However, there are error messages in the output of the plotrgascan MATLAB script; evidiently there are some negative/bogus values in the output. 

I'll look into it more tomorrow.

  12114   Tue May 10 03:44:59 2016 ericqUpdateLSCRelocked

ALSX noise is solidly within past acceptable performance levels. The DRFPMI was locked on four out of six attempts. 

Some housekeeping was done:

  • PMC aligned
  • Static alignment voltages of X end PZT mirrors offloaded by turning mount screws
  • Rough comissioning of AUX X dither alignment
  • Locking scripts reverted to AUX X Innolight voltage/temperature sign convention

The recombination of the QPD signals to common / differential is imperfect, and limited how well we could keep the interferometer aligned, since the QPD at X has changed. This needs some daytime work. 

Some sensing matrix measurements were made, to be meditated upon for how to 1F the DRMI.

Other to-dos:

  • Bandpass + notch combo for green refl PDs
  • SRCL, and to a lesser extant, MICH feedforward subtraction (See DARM vs. other length DOF coherence plot below)
  • Fiber couple AUX X light
  • Make IFO work good


As an aside, Gautam and I noticed numerous green beams coming from inside the vacuum system onto the PSL table. They exist only when green is locked to the arms. Some of them come out at very non-level angles and shine in many places. This doesn't make me feel very happy; I suppose we've been living with it for some time. 

Attachment 1: 2016-05-10_DARMcoherence.pdf
2016-05-10_DARMcoherence.pdf
  12113   Sun May 8 08:39:21 2016 ranaUpdateLSCGreen PDH demod lowpass

Indeed. This is why the LSC PDs have a 2f notch in addition to the 1f resonance. In recent versions, we also put a 2f notch in the feedback of the preamp which comes after the diode but before the mixer. The overall 1f to 2f ratio that we get is in the 50-60 dB region. I don't think we have to go that far with this thing; having a double LC already seems like it should be pretty good, or we could have a single LC bandpass with a 2f notch all in one Pomona box.

  12112   Sat May 7 09:40:40 2016 ericqUpdateLSCGreen PDH demod lowpass

As I was looking at filter designs, it seemed difficult to get 40dB of supression at 2F with a bandpass without going to a pretty high order, which would mean a fair number of lossy inductors.

I'll keep working on it. Maybe we don't need 40dB...

  12111   Fri May 6 19:08:52 2016 ranaUpdateLSCGreen PDH demod lowpass

Seems weird to design a PD lowpass with a corner at the modulation frequency. Recall what our strategy is with the other photodetectors we use for PDH servos: bandpass, not low-pass, and the band has to be wide enough to not effect the phase of the servo.

  12110   Fri May 6 16:42:12 2016 ericqUpdateLSCGreen PDH demod lowpass

I've build the filter, and it seems to have the desired TF shape.

I also re-purposed the 70k lowass to a ~120k lowpass by changing the 68nF caps to 22nF caps, since we still want some post-mixer rolloff. 

However, putting the ELPF in the chain caused some weird shapes in the OLG. I still need to get to the bottom of it. However, just with the post-mixer LPF modification, here's what the OLG looks like:

As Rana surmises, we definitely still add a boost and maintain a 10k UGF. I still need to look into the state of the remote boost....

  12109   Thu May 5 21:28:44 2016 gautamUpdateendtable upgradeInnolight PZT capacitance

I suggested in an earlier elog that after the repair of the NPRO, the PZT capacitance may have changed dramatically. This seems unlikely - I measured the PZT capacitance with the BK Precision LCR meter and found it to be 2.62 nF, which is in excellent agreement with the numbers from elogs 3640 and 4354 - but this makes me wonder how the old setup ever worked. If the PZT capacitance were indeed that value, then for the Pomona box design in elog 4354, and assuming the PM at ~216kHz which was the old modulation frequency was ~30rad/V as suggested by the data in this elog, we would have had a modulation depth of 0.75 if the Function Generator were set to output a Signal at 2Vpp (2Vpp * 0.5 * 0.05 * 30rad/V = 1.5rad pp)! Am I missing something here?

Instead of using an attenuator, we could instead change the capacitor in the pomona box from 47pF mica to 5pF mica to realize a modulation depth of ~0.2 at the new modulation frequency of 231.25 kHz. In any case, as elog 4354 suggests, the phase introduced by this high-pass filter is non-zero at the modulation frequency, so we may also want to install an all-pass filter which will allow us to control the demodulation phase. This should be easy enough to implement with an Op27 and passive components we have in hand...

 

  12108   Thu May 5 14:05:01 2016 ranaUpdateendtable upgradeALS status update

All seems very fishy. Its not good to put attenuators and filters in nilly-willy.

  1. Once the post-PD bandpass has been designed and constructed, you should be able to use whatever PD gain setting gives you the best SNR. There's no need to use more PD gain than necessary; it just reduces the PD bandwidth. What is the input referred current noise of the PD at the different gain settings?
  2. The open loop mixer output *should* be very large. It should be reduced to mV only when the loop is closed.
  3. The better way to estimate the modulation depth is to lock the arm on red as usual and then scan the EX laser and look at the green transmission. The FSR is 3.7 MHz, so the SBs should show up well in a narrow scan around the carrier.
  4. I guess its going to be tough to impedance match the splitter box to the NPRO PZT, since its impedance is all over the place at 200-300 kHz, but you could put a 50 Ohm in-line terminator in there somewhere?
  5. The Bode plot seems to indicate that we could easily get a 10 kHz UGF and then switch on a Boost. Is the remote Boost switch disabled or always ON? I am suspicious of the plot and think that the coarse trace is probably missing some sharp resonances which will sneakily bite you.
  12107   Thu May 5 14:03:52 2016 ericqUpdateLSCFurther Aux X PDH tweaks

This morning I poked around with the green layout a bit. I found that the iris immediately preceding the viewport was clipping the ingoing green beam too much, opening it up allowed for better coupling to the arm. I also tweaked the positions of the mode matching lenses and did some alignment, and have since been able to achieve GTRX values of around 0.5.

I also removed the 20db attenuator after the mixer, and turned the servo gain way down and was able to lock easily. I then adjusted the gain while measuring the CLG, and set it where the maximum gain peaking was 6dB, which worked out to be a UGF of around 8kHz. On the input monitor, the PDH horn-to-horn voltage going into the VGA is 2.44V, which shouldn't saturate the G=4 preamp stage of the AD8336, which seems ok.

The ALS sensitivity is now approaching the good nominal state:

There remains some things to be done, including comprehensive dumping of all beams at the end table (especially the reflections off of the viewport) and the new filters to replace the current post-mixer LPF, but things look pretty good.

Attachment 1: 2016-05-05_newals.pdf
2016-05-05_newals.pdf
  12106   Thu May 5 04:05:03 2016 ericqUpdateLSCAux X PDH checks

We took an OLG measurement of the green PDH loop. It seems consistent with past measurements. I've added a trace for the the post-mixer lowpass, to show its contribution to the phase loss. (EDIT: updated with measured LPF TF)

I used this measured OLG and the datasheet laser PZT conversion factor to calibrate the control signal monitor into the AUX laser frequency noise, it looks consistent with the frequency noise measured via the PSL PLL (300 Hz/rtHz @ 100Hz). Above a few tens of kHz, the control signal measurement is all analyzer noise floor, due to the fourth order 70kHz lowpass after the mixer (the peaks change height significantly depending on the analyzer input range, so I don't think they're on the laser). Gautam will follow up with more detailed measurements of both the error and control signals as he noisebudgets, this was just intended as a quick consistency check.

  12105   Thu May 5 03:05:37 2016 gautamUpdateendtable upgradeALS status update

[ericQ, gautam]

Today we spent some time looking into the PDH situation at the X end. A summary of our findings.

  1. There is something that I don't understand with regards to the modulation signal being sent to the laser PZT via the sum+HPF pomona box - it used to be that with 2Vpp signal from the function generator, we got ~5mVpp signal at the PZT, which with the old specs resulted in a modulation of ~0.12rad. Now, however, I found that there was a need to place a 20dB attenuator after the splitter from the function generator in order to realize a modulation depth of ~0.25 (which is what we aim for, measured by locking to the TEM00 modes of the carrier and sidebands and comparing the ratio of powers). It could be that the PZT capacitance has changed dramatically after the repair. Nevertheless, I still cant reconcile the numbers. We measured the transfer function from the LO input of the pomona box to the output with the PZT connected, and figure there should be ~70dB of attentuation (with the 20dB additional attenuator in place). But this means 1Vpp*0.0003*70rad/V = 0.02rad which is an order of magnitude away from what the ratio of powers suggest. Maybe the measurement technique was not valid. In any case, this setup appears to work, and I'm also able to send +7dBm to the mixer which is what it wants (function generator output is 3Vpp).
  2. In addition to the above, I found that the demodulated error signal had a peak-to-peak of a few volts. But the PDH servo is designed to have tens of mV at the input. Hence, it was necessary to turn down the gain of the REFL PD to 10dB and add a 20dB attenuator between mixer output and servo input.
  3. While Johannes and I were investigating this earlier in the afternoon, we found that the waveform going to the laser PZT was weirdly distorted (still kind of sinusoidal in shape, but more rounded, I will put up a picture shortly). This may not be the biggest problem, but perhaps there is a better way to pipe the LO signal to the PZT and mixer than what is currently done.
  4. We then looked at loop transfer function and spectrum of the control signal. Plots to follow. They look okay.
  5. I measured the green power coming onto the PSL table. It is ~400uW. After optimizing alignment, the green transmission is ~0.4 according to whatever old normalization we are using.
  6. We then recovered the X green beatnote and looked at the ALS noise spectrum. Beatnote amplitude at the beat PD is ~ -27dBm. The coherence in the region of a few hundred Hz suggests that some improvements can be made to the PDH situation (the gain of the PDH servo is maxed out at the X end at the moment...). But the bottom line is this is probably good enough to get back to locking...
Attachment 1: ALS_noiseSpec_5May2016_2.pdf
ALS_noiseSpec_5May2016_2.pdf
Attachment 2: Coherence_5May2016.pdf
Coherence_5May2016.pdf
Attachment 3: image.jpeg
image.jpeg
  12104   Mon May 2 19:14:18 2016 gautamUpdateendtable upgradeOptical layout almost complete

With Steve's help, I installed the Oplev earlier today. I adjusted the positions of the two lenses until I deemed the spot size on the QPD satisfactory by eye. As a quick check, I verified using the DTT template that the UGF is ~5Hz for both pitch and yaw. There is ~300uW of power incident on the QPD (out of ~2mW from the HeNe). In terms of ADC counts, this is ~13,000 counts which is about what we had prior to taking the endtable apart. There are a couple of spots from reflections off the black glass plate in the vacuum chamber, but in general, I think the overall setup is acceptable.

This completes the bulk of the optical layout. The only bits remaining are to couple the IR into the fiber and to install a power monitoring PD. Pictures to follow shortly. 

Now that the layout is complete, it remains to optimize various things. My immediate plan is to do the following:

  1. Maximize green transmission by tweaking alignment. I should also do a quick check using mirror specs to see that the measured transmitted green power compares favourably to what is expected.
  2. Check the green PDH loop transfer function at the X end - this will allow me to set the gain on the uPDH box systematically.
  3. Re-establish green beats, check noise performance.
  4. There are possibly multiple beam dumps that have to be installed. For now, I've made sure that no high power IR beams are incident on the enclosure. But there are a couple of red and green beams that have to be accounted for.

I will also need to upload the layout drawing to reflect the layout finally implemented.


Not directly related:

The ETMx oplev servo is now on. I then wanted to see if I could lock both arms to IR. I've managed to do this successfully - BUT I think there is something wrong with the X arm dither alignment servo. By manually tweaking the alignment sliders on the IFOalign MEDM screen, I can get the IR transmission up to ~0.95. But when I run the dither, it drives the transmission back down to ~0.6, where it plateaus. I will need to investigate further. 

 

GV Edit: There was some confusion while aligning the Oplev input beam as to how the wedge of the ETM is oriented. We believe the wedge is horizontal, but its orientation (i.e. thicker side on the right or left?) was still ambiguous. I've made a roughly-to-scale sketch (attachment #1) of what I think is the correct orientation - which turns out to be in the opposite sense of the schematic pinned up in the office area.. Does this make sense? Is there some schematic/drawing where the wedge orientation is explicitly indicated? My search of the elog/wiki did not yield any..

Attachment 1: ETMX_wedge.pdf
ETMX_wedge.pdf
  12103   Mon May 2 17:11:55 2016 ranaUpdateCOCRC folding mirrors

Antonio/Gautam are now developing a more up to date Finesse model of our recycling cavities to see what we can have there before our power recycling gain or cavity geometric stability is compromised. Expect that we will here a progress report on the model on Wednesday.

Some thoughts:

  1. RC folding mirrors need to be dichroic to allow green beams to get out.
  2. We should look at the specs Jamie used to get the RC folding mirrors last time and figure out what went wrong / what specs to change.
  3. T_1064 < 100 ppm. Hopefully < 50 ppm.
  4. On the AR side, we mainly want low AR for green, but nothing special for 1064, since that's taken care of by the HR.
  5. How much should we wedge these things?
  6. Should the wedge be horizontal?
  7. Can we get someone in Downs to update the optical layout?
  8. What microroughness do we need?
  9. The mirrors must be flat, with the  500 m < RoC < 100 km. Part of the Finesse modeling is to figure out what happens if the RoC is in the 300 - 1000 m range. Better stability?
  12102   Mon May 2 17:06:58 2016 ranaSummaryCOCG&H optics to Fullerton/HWS for anneal testing

Steve sent 4 of our 1" diameter G&H HR mirrors to Josh Smith at Fullerton for scatter testing. Attached photo is our total stock before sending.

Attachment 1: 20160427_182305.jpg
20160427_182305.jpg
  12101   Fri Apr 29 16:13:36 2016 ericqUpdateLSCGreen PDH demod lowpass

We can get as much, if not more, attenuation of the 1F line in the mixer output that we get from the post-mixer LPF from using the following passive filter between the PD and mixer RF input:

There should still be some kind of LPF after the mixer, but I haven't yet determined what it should be; this will determine how much phase the PDH loop wins. At most, this should win around 25 degrees at 10kHz.


The filter was designed by referencing the "Handbook of Filter Synthesis" by Zverev, looking for an elliptic filter for matched source and load impedences, 40dB min attenuation in the stopband, a stopband frequency that starts at twice the corner frequency, and minimizing the VSWR between the PD and filter in the passband.

In terms of the tables in the book, this means: n=5, rho=2%, theta=30deg, K**2 = 1.0. The dimensionless component values were scaled by the corner frequency of 200kHz, and reference impedence of 50 Ohm. (The corner is a little lower than the real modulation frequency, since the nonzero resistance of the inductors pushes the frequency up a bit)

The ideal capactior values do not correspond to things we have in hand, so I checked our stock and chose the closest value to each one.Unsurprisingly, due to these component substitutions, and the fact that the coilcraft inductors have a resistance of about 7 Ohms, the predicted TF of the realizable filter does not match the design filter exactly. However, the predicition still looks like it will meet the requirement of 40dB of supression of the 2F line in the PD signal. (Since we have tunable inductors, I've used the ideal inductor values in generating the TF. In practice I'll inspect the TF while I tune them)

  Desired Realizable
C1 8.28 nF 10 nF
C2 1.39 nF 1.5 nF
C3 19.6 nF 22 nF
C4 4.22 nF 4.7 nF
C5 6.08 nF 6.8 nF
L2 43.1 nH 32-48 nH + 7 Ohm
L4 34.4 nH 32-48 nH + 7 Ohm

[In this TF plot, I've multiplied the real response by 2 to account for the voltage division that occurs with ideal 50 Ohm impedance matching, to make 0dB the reference for proper matching]

The filter's phase delay at the modulation frequency is just about 180, which as a time delay of 5usec works out to 9 degrees of phase loss at 10kHz in the PDH loop. According to some old measurements, the current LPF costs something like 35 degrees at 10k, so this wins at most around 25 degrees, depedent on what LPF we put after the mixer.

LISO source both traces is attached!

Attachment 3: elp_liso.zip
  12100   Fri Apr 29 16:05:23 2016 gautamUpdateendtable upgradeCleaning ETMX vacuum dirty window

After a second round of F.C. application, I think the window is clean enough and there are no residual F.C. pieces anywhere near the central parts of the window (indeed I think we got most of it off). So I am going to go ahead and install the Oplev. 

Quote:

It looks very promising.

 

 

Attachment 1: IMG_0755.JPG
IMG_0755.JPG
  12099   Fri Apr 29 00:55:46 2016 gautamUpdateendtable upgradegreen PDH locked to Xarm

Using the modulation frequency suggested here, I hooked up the PDH setup at the X-end and succeeded in locking the green to the X arm. I then rotated the HWP after the green Faraday to maximize TRX output, which after a cursory alignment optimization is ~0.2 (I believe we were used to seeing ~0.3 before the end laser went wonky). Obviously much optimization/characterization remains to be done. But for tonight, I am closing the PSL and EX laser shutters and applying first contact to the window once more courtesy more PEEK from Koji's lab in W Bridge. Once this is taken care of, I can install the Oplev tomorrow, and then set about optimizing various things in a systematic way.. MC autolocker has also been disabled...

Side note: for the IR Transmon QPD, we'd like a post that is ~0.75" taller given the difference in beam height from the arm cavity and on the endtable. I will put together a drawing for Steve tomorrow..

  12098   Thu Apr 28 18:53:05 2016 ranaUpdateLSCGreen PDH demod lowpass

OK - but give us a circuit diagram and the expected before/after loop plots. Got to make sure we keep the right impedance from PD to mixer. Some of the Thorlabs PDs have a 50 Ohm instead of 0 Ohm source impedance. Maybe you can try it out now since the green arm is ready.

  12097   Thu Apr 28 15:23:11 2016 ericqUpdateLSCGreen PDH demod lowpass

The 2F product out of the mixer is a natural concern when demodulating. However, I think this isn't so big of a deal in our green PDH servos; 420kHz isn't so high of a frequency that the servo amplifiers are bandwidth or slew-rate limited. Furthermore, the amplitude of this line is supressed by the loop somewhat, since it arises from the same field product that the loop is acting on. Measuring the Y end mixer output with a high impedance probe and the AG4395 shows it to be something like -50dBm. 

In fact, the main thing that the pomona LPFs are accomplishing right now is filtering the 1F content of the mixer output that arises from the second order sideband creating a signal at 2F, and beating with the LO at (2F-1F)=1F. This line is something like -30dBm (5mVrms) at the mixer output; I can reproduce this amplitude with a back-of-the envelope calculation using a modulation depth of 0.3, 8V out of the PD at DC when unlocked, the mixer datasheet, and the nominal cavity parameters. 

The nice thing about this is that we don't need to filter this after the mixer, we can use a [bandpass/lowpass/notch] filter before the mixer (as is done in the LSC demod boards) to filter out the 2F (420kHz) content of the PD signal, which will only introduce some small amount of linear time delay to the PDH loop, instead of the wicked phase loss from the current post-mixer LPF. We can then replace that 70kHz filter with something of lower order or higher corner frequency to win a good deal of phase in the PDH loop. 

  12096   Thu Apr 28 08:49:47 2016 SteveUpdatePEMpossible noise sources schedule

Building:         Campus (see attached Map)        

       

Date:              Manhole 1 - May 3 through May 5

Manhole 2 – May 6 through May 10

 Manhole 2 - May 16 through May 19

Manhole 3 – May 11 through May 19           

          

Time:              Noise:  7:00 a.m. To 5:00 p.m.

                  Access: 24 Hours a day

           

Interruption:      Noise/Vehicular & Pedestrian Access

                  Storm Drain Manholes

         

*In order to repair 3 manholes associated with a large storm drain that runs north-south through the campus, work will take place at the

3 manholes shown on the map. This work will interrupt vehicular and pedestrian access on the paths adjacent to the manholes. Though the work at Manholes 1 and 2 will allow vehicular and pedestrian access around the manholes, the work at Manhole 3 will completely block the driveway running south from the southeast corner of Parking Lot 11. Noise will also be created by the repair

 

 

Attachment 1: Campus_B&W_Map-2.pdf
Campus_B&W_Map-2.pdf
  12095   Thu Apr 28 00:41:08 2016 gautamUpdateendtable upgrademore progress - Transmon PD installed

The IR Transmon system is almost completely laid out, only the QPD remains to be installed. Some notes:

  1. The "problem" with excessive green power reflected from the harmonic separator has been resolved. It is just very sensitive to the angle of incidence. In the present configuration, there is ~10uW of green power reflected from either side, which shouldn't be too worrisome. But this light needs to be dumped. Given the tiny amount, I think a black glass + sticky tape solution is best suited, given the space constraints. This does not reach the Transmon PDs because there is a filter in the path that is transmissive to IR only. 
  2. I aligned the transmitted beam onto the Thorlabs PD, and reconnected the signal BNC cable (the existing cable wasn't long enough so I had to use a barrel connector and a short extension cable). I then reverted the LSC trigger for the X arm back to TRX DC and also recompiled c1ass to revert to TRX for the dither alignment. At the moment, both arms are stably locked, although the X arm transmission is saturated at ~0.7 after running the dither alignment. I'm not sure if this is just a normalization issue given the new beam path or if there is something else going on. Further investigations tomorrow.
  3. It remains to dump some of the unwanted green light from the addition of the harmonic separator...
  4. We may want to redesign some (or all) of the Transmon path - the lens currently in use seems to have been chosen arbitrarily. Moreover, it is quite stubbornly dirty, there are some markings which persist after repeated first contact cleaning...

I feel like once the above are resolved, the next step would be to PDH lock the green to the arm and see what sort of transmission we get on the PSL table. It may be the polarization or just alignment, but for some reason, the transmitted green light from the X arm is showing up at GTRY now (up to 0.5, which is the level we are used to when the Y arm has green locked!). So a rough plan of action:

  1. Install transmon QPD
  2. PDH lock green to X arm
  3. Fix the window situation - as Steve mentioned in an earlier elog, the F.C. cleaning seems to have worked well, but a little remains stuck on the window (though away from where any laser radiation is incident). This is resolved easily enough if we apply one more layer of F.C., but the bottle-neck right now is we are out of PEEK which is what we use to remove the F.C. once dried. Steve thinks a fresh stock should be here in the next couple of days...
  4. Once 3 is resolved, we can go ahead and install the Oplev.
  5. Which leaves the lst subsystem, coupling to the fiber and a power monitor for the NPRO. I have resolved to do both these using the 1% transmitted beam after the beamsplitter immediately after the NPRO rather than pick off at the harmonic separator after the doubling oven. I need to do the mode-matching calculation for coupling into the fiber and also adjust the collimating lens...
  6. Clean-up: make sure cables are tied down, strain-relieved and hooked up to whatever they are supposed to be hooked up to...
  12094   Wed Apr 27 15:04:47 2016 SteveUpdateendtable upgradeCleaning ETMX vacuum dirty window

It looks very promising.

 

Attachment 1: 1cETMX-Tcmp.jpg
1cETMX-Tcmp.jpg
  12093   Wed Apr 27 14:06:31 2016 ranaSummaryGeneralmeeting notes
  1. Gautam will get help from Johannes and finish EX table by Monday.
  2. Steve will spend a day this week with Johannes on Green Monster bakeout.
  3. Q to analyzed green PDH servo and design demod low pass. Should we use the double LC notches to notch the 2f product? What's the demod filter attenuation requirement?
  4. Koji will make a drawing of the ruby suspension standoff prism and post into the elog so that Steve can get some quotes next week.
  5. Rana to implement 40m configuration in FOGprime17 and analyze RoC matching of ETMs. Get Antonio's help to analyzed SRC stability. Maybe use PyKat and Finesse since Antonio knows that stuff.
  6. Give OCXO boxes to WB refcav people. Rana get Rich to make another couple of boxes for 40m PMC, FSS, IMC.
  7. Rana/Koji get EKG to make specs and procure some new folding mirrors for the PRC/SRC. Make them a bit concave and dichroic.
  12092   Wed Apr 27 09:45:56 2016 ranaUpdateSUSspare SOS tower

Bah, we need ruby slippers for all future suspensions. Prism with curved backside and smooth grooves.

No aluminum, no cry.

Quote:

Earth quake stops need viton tips.

Wirestandoffs are still aluminum.

 

  12091   Wed Apr 27 09:05:10 2016 SteveUpdateGeneralAP viewport

                   Sad situation

    The anti-symmetric port

spider webs fly in the wind

Attachment 1: APg.jpg
APg.jpg
Attachment 2: AP.jpg
AP.jpg
Attachment 3: APspiderWebs.jpg
APspiderWebs.jpg
  12090   Tue Apr 26 23:19:42 2016 gautamUpdateendtable upgradeGreen aligned to arm - high order mode flashes seen

Attachment #1

Layout as of today. Most of the green path is done. The Green REFL PD + PZT mirrors have not been hooked up to their respective power sources yet (I wonder if it's okay to start laying cables through the feedthroughs on either end of the table already, or if we want to put whatever it is that makes it airtight eventually in first?). A rough power budget has been included (with no harmonic separator just before the window), though some optimization can be done once the table is completely repopulated.

Attachment #2

A zoomed-in version of the REFL path.

Some general notes:

  1. I've tried to use the custom 3/4" O.D. posts + baseplate arrangement wherever possible (only 1 steering mirror is on a 1" post clamped with a fork to the table because of space constraints). Where the baseplates could not be bolted onto the table directly, I've used Newport SS Dogs to do the job.
  2. I checked for continuity between the PZT outer case and the table top with a multimeter, and found none. So I chose to leave the Thorlabs baseplates in place. For the REFL PD, I've used an insulating baseplate given to me by Steve.
  3. I've used some custom length 3/4" O.D. posts to get the beam up to the right height (~4.75") just before sending the green beam in. The beam height is 4" elsewhere.
  4. I was playing around with positioning the harmonic separator immediately before the vacuum chamber window - I found that there is a substantial amount of green light that is reflected, though there doesn't seem to be any IR leaking through. The mirror was labelled Y1-1037-45P, which is a code for CVI mirrors, though I believe it is a LaserOptik product and that we have a couple of other such mirrors in the optics cabinet - though they are all 1". This document suggests that from the back side, there should be <0.1% reflection of green while on the front side it should be < 3%. I will have to hunt a little more for the specs, and measure the powers to see if they match the previously quoted numbers. In any case, I'll have to think of how to separate the (unwanted) reflected green and the transmitted IR from the cavity in the IR transmon path.
  5. There are some minor changes to the planned layout posted here - I will update these in due course once the Transmon path and Oplev have been set up.

I am closing the PSL shutter and the EX laser shutters for the night as I have applied a layer of first contact to the window for cleaning purposes, and we don't want any laser light incident on it. It may be that the window is so dirty that we may need multiple F.C. cleaning rounds, we will see how the window looks tomorrow...

 

Attachment 1: IMG_2219.JPG
IMG_2219.JPG
Attachment 2: IMG_2220.JPG
IMG_2220.JPG
  12089   Tue Apr 26 15:22:35 2016 SteveUpdateendtable upgradeCleaning ETMX vacuum dirty window

Gautom is progressing with the layout nicely. The X-arm transmission window have not seen cleaning for decades. This should be the time to do it. Here is picture of dirtiness.

It is not that simple... How much effort should we put in it? The hole table with 1W inno laser plus... set up now about ~500 lbs  We can pull it off carefully, but it is not risk free.

We should look at our other signal port windows! Gautom's long reach able him to do the first contact cleaning without moving anything. It is great!
 

Attachment 1: ETMX-Tvp.jpg
ETMX-Tvp.jpg
Attachment 2: ETMX-TvpDetail.jpg
ETMX-TvpDetail.jpg
  12088   Mon Apr 25 11:07:06 2016 SteveUpdateSUSspare SOS tower

Earth quake stops need viton tips.

Wirestandoffs are still aluminum.

Attachment 1: ETMXreplacment.jpg
ETMXreplacment.jpg
  12087   Fri Apr 22 13:58:13 2016 SteveUpdatePEMleaky roof is fixed

Dan sealed the leak today.

 

Attachment 1: leakyRoof_(2).jpg
leakyRoof_(2).jpg
  12086   Thu Apr 21 15:12:38 2016 SteveUpdateVACRGA is not working

 

Quote:

Steve pointed out that in the aftermath of the Nitrogen running out a couple of times last week, the RGA had shut itself off thinking that there was a leak and so it was not performing the scheduled scans once a day. So the data files from the scheduled scans were empty in the /opt/rtcds/caltech/c1/scripts/RGA/logs directory. The wiki page for getting it up and running again is up-to-date, but the script RGAset.py did not exist on the c0rga machine, which the RGA is communicating with via serial port. I copied over the script RGAset.py from rossa to c0rga and ran the script on that machine - but the error flags it returned were not all 0 (indicating some error according to the manual) - so I edited the script to send just the initialize command ('IN0') and commented out the other commands, after which I got error flags which were all 0. After this, I ran a manual scan using 'RGAlogger.py', and it appears that the RGA is now able to take scans again - I'm attaching a plot of the scan results. We've saved this scan as a reference to compare against after a few days. 

Our last RGA scan is from February 14, 2016  We had a power outage on the 15th

Gautom has not succeded  reseting it. The old c0rga computer looks dead. Q may resurrect it, if he can?

  12085   Thu Apr 21 14:25:52 2016 gautamUpdateendtable upgradeGreen light recovered

I've made progress on the new layout up to the doubling oven. After doing the coarse alignment with the diode current to the NPRO at ~1A, I turned it back up to the nominal 2A. I then rotated the HWP before the IR Faraday such that only ~470mW of IR power is going into the doubler (the rest is being dumped on razor beam dumps). After tuning the alignment of the IR into the doubling oven using the steering mirror + 4 axis translation stage on which the doubling oven is mounted, I get ~3.2mW of green after the harmonic separator and a HR mirror for green. The mode looks pretty good to the eye (see attachment #1), and the conversion efficiency is ~1.45%/W - which is somewhat less than the expected 2%/W but in the ballpark. It may be that some fine tweaking of the alignment + polarization while monitoring the green power can improve the situation a little bit (I think it may go up to ~4mW, which would be pretty close to 2%/W conversion efficiency). The harmonic separator also seems to be reflecting quite a bit of green light along with IR (see attachment #2) - so I'm not sure how much of a correction that introduces to the conversion efficiency. 

While doing the alignment, I noticed that some amount of IR light is actually transmitted through the HR mirrors. With ~500mW of incident light at ~45 degrees, this transmitted light amounts to ~2mW. Turns out that this is also polarization dependant (see attachment #3) - for S polarized light, as at the first two steering mirrors after the NPRO, there is no transmitted light, while for P-polarized light, which is what we want for the doubling crystal, the amount transmitted is ~0.5%. The point is, I think the measured levels are consistent with the CVI datasheet. We just have to take care find all these stray beams and dump them.

I will try and optimize the amount of green power we can get out of the doubler a little more (but anyway 3mW should still be plenty for ALS). Once I'm happy with that, I will proceed with laying out the optics for mode-matching the green to the arm.

Attachment 1: IMG_6567.JPG
IMG_6567.JPG
Attachment 2: IMG_6568.JPG
IMG_6568.JPG
Attachment 3: CVI_reflectivity.jpeg
CVI_reflectivity.jpeg
ELOG V3.1.3-